# DamageCAT: A Deep Learning Transformer Framework for Typology-Based Post-Disaster Building Damage Categorization

### v1

@@ 4.3 Challenges and limitations
Despite the promising results, this work faces several key limitations that constrain its immediate scalability and generalizability. First, there is limited access to consistent, high-resolution pre- and post-disaster imagery pairs; obtaining uniform spatial resolution and temporal alignment can be challenging, particularly in large-scale or multi-hazard contexts. Additionally, manual annotation remains time-intensive, especially for subtle or nuanced damage categories (e.g., partial roof damage vs. partial structural collapse), which can slow dataset expansion and increase labeling costs.
From a methodological perspective, relying solely on vertical (nadir) satellite imagery can obscure certain facade-level or internal damages, leading to potential misclassifications or omissions. Moreover, our dataset exhibits significant class imbalance, most notably in the total structural collapse category (representing only 0.47% of annotated buildings), which can affect model robustness despite mitigation strategies like weighted losses. The focus on a single disaster event (Hurricane Ida) and one geographic region (Louisiana) further narrows the framework’s initial scope. As Adriano et al. [3] cautions, damage assessment methods tailored to specific locales may have limited transferability to different hazards or global settings.
Finally, computational requirements—arising from the transformer-based architecture—may hinder real-time deployment in resource-constrained environments, such as remote field operations or developing regions. In line with Gholami et al. [1], inference speed can be more critical than absolute accuracy during crises, underscoring the need for further model optimization (e.g., model distillation, quantization) to meet operational demands. Addressing these limitations through multi-modal data integration, lighter model architectures, and extensive cross-regional validation represents an important avenue for future research.
@@ 4.4 Future research directions
Future research should focus on developing balanced datasets across damage categories, disaster types, and geographic regions to enhance model generalizability. Integrating multiple data sources (e.g., oblique aerial photography, street-view imagery) to provide complementary perspectives, as suggested by Braik and Koliou [36] and Xue et al. [39], and incorporating multiple post-disaster time points to track recovery progression are also important. Furthermore, developing interpretable models that highlight visual features contributing to classifications and optimizing models for edge devices with limited computational resources are key areas for advancement. These initiatives require interdisciplinary collaboration to develop systems that provide actionable intelligence for disaster response.
@@ 5 Conclusion
This study introduced DamageCAT, a novel framework for typology-based building damage assessment that transcends the limitations of purely severity-focused methods. By proposing the BD-TypoSAT dataset—a human-annotated resource containing four distinct damage categories (partial roof damage, total roof damage, partial structural collapse, and total structural collapse)—we address the pressing need for descriptive, real-world training data that captures the multifaceted nature of post-disaster structural impairments. Concurrently, we developed a hierarchical U-Net-based transformer architecture capable of analyzing pre- and post-disaster satellite imagery to produce accurate and highly detailed damage classification maps. Through comprehensive experiments, our model achieved robust performance metrics (IoU: 0.7921, F1: 0.8835), demonstrating its ability to detect both common and rare damage types, thereby offering a powerful new tool for emergency responders.
Beyond the quantitative success, the novelty of DamageCAT lies in its paradigm shift from broad severity categories to typology-based assessment, which provides fine-grained insights essential for targeted resource allocation and actionable decision-making. By pinpointing whether a building sustains roof damage versus structural collapse, the framework fosters improved prioritization of personnel, materials, and financial support. The model’s reliance on pre-post disaster change detection also reduces false positives by filtering out pre-existing building defects, further enhancing its practical utility in time-critical scenarios.
Despite these promising results, the study faces a few key limitations. First, the dataset’s imbalance (particularly for total structural collapse) reflects the real-world rarity of certain extreme damage types, yet indicates a need for expanded, balanced datasets to fully optimize performance. Similarly, the approach currently relies on vertical satellite imagery, which may obscure critical facade-level or internal damage features. Finally, the focus on a single disaster event (Hurricane Ida) and region (New Orleans, Louisiana) limits initial generalizability, and the high computational demands of the transformer-based design may pose challenges for real-time deployment in resource-constrained contexts.
Looking ahead, future research should collect and annotate more balanced data across multiple hazards and geographies, potentially integrating multi-modal sources (e.g., oblique aerial imagery, street-view photographs, LiDAR) to broaden the model’s detection capabilities. Addressing real-time inference constraints will be critical for on-site deployment, particularly by exploring model optimization techniques (e.g., pruning, and quantization). Additionally, the development of interpretable AI methods could boost user trust and offer transparency into the visual cues the model associates with specific damage types. Such innovations, combined with cross-regional validations and in-field performance tests, will continue to refine the typology-based assessment paradigm and ensure the widespread adoption of advanced building damage classification in post-disaster management.
Future research should prioritize real-time inference for emergency deployment, integration of pre-disaster building characteristics, and field validations in diverse contexts. Additionally, using interpretable AI techniques could enhance performance and user trust by revealing the visual features influencing classification.
@@
Data and Code availability The data and code that support the findings of this study are available from the corresponding authors upon request.

### v2

@@ 5.3 Challenges and limitations
Despite promising results, several key limitations constrain the framework’s immediate scalability and generalizability. These challenges span data acquisition, methodological constraints, and practical deployment considerations.
Data acquisition and costs represent the primary barrier to widespread adoption. While building-level damage assessment requires sub-meter resolution imagery ($\sim$0.5 m/pixel), most freely available sources like Sentinel-1/2 and Landsat provide only 10–30 m resolution—adequate for flood mapping but insufficient for distinguishing roof damage from structural collapse. Commercial alternatives are costly: as shown in Table 5, high-resolution imagery ranges from \$2.25/km2 for 3 m archives to \$22.50/km2 for 30 cm data, with ultra-high-resolution SAR costing \$675–\$3,250 per 5 km $\times$ 5 km scene. Even when emergency providers like Maxar Open Data or Planet Labs donate post-disaster imagery, comprehensive pre-disaster coverage typically requires commercial acquisition. Manual annotation compounds these costs, particularly for subtle damage distinctions that demand expert interpretation.
Methodological limitations stem from relying solely on nadir satellite imagery, which can miss facade-level or internal damage visible only from oblique angles. Severe class imbalance—with total structural collapse representing just 0.47% of annotated buildings—affects model robustness despite weighted loss mitigation. The framework’s training on a single event (Hurricane Ida) in one region (New Orleans, Louisiana) also limits generalizability, as damage patterns vary significantly across hazard types and building typologies [3].
Computational requirements from the transformer architecture may hinder real-time deployment in resource-constrained environments. During crises, inference speed often matters more than marginal accuracy gains [1], suggesting the need for model optimization techniques like distillation or quantization.
Addressing these limitations through multi-modal data integration, lighter model architectures, and extensive cross-regional validation represents an important avenue for future research.
Table 5: Indicative costs and access models for satellite imagery in disaster response.
Provider / Product Native resolution Typical cost or licence Source
PlanetScope (archive) 3 m \$2.25/km2 (min 250 km2) [47]
Pléiades Neo (archive) 30 cm \$22.50/km2 (min 25 km2) [48]
Maxar Open Data 30–50 cm Free for sudden-onset disasters [49]
Planet Labs 3–5 m Free on request for accredited responders [50]
Umbra SAR tasking 0.25–1 m radar \$675–\$3,250 per 5 km $\times$ 5 km scene [51]
Copernicus EMS 0.5–10 m (mixed) Free maps; imagery cost covered by EU [52]
Sentinel 1/2 10–20 m optical/radar Free open access [53]
Landsat archive 15–30 m Free open access [54]
International Charter Mixed sensors Free emergency tasking (authorised users) [55]
@@ 5.4 Future research directions
Several research directions emerge from this work’s limitations and findings. Dataset expansion represents the most critical need, requiring balanced representation across damage categories, disaster types, and geographic regions to improve model generalizability. Current class imbalances and regional specificity limit broader applicability, suggesting the need for coordinated data collection efforts across multiple hazard events.
Multi-modal data integration offers significant potential for improving classification accuracy. Incorporating oblique aerial photography and street-view imagery could provide complementary perspectives on building damage that nadir satellite imagery cannot capture [41, 44]. Additionally, integrating multiple post-disaster time points would enable tracking of recovery progression and damage evolution over time.
Model interpretability and optimization represent important technical directions. Developing interpretable models that highlight specific visual features contributing to damage classifications would enhance trust and adoption among emergency responders. Simultaneously, optimizing architectures for edge devices with limited computational resources is essential for field deployment in resource-constrained environments.
Operational integration requires interdisciplinary collaboration between computer scientists, disaster researchers, and emergency management practitioners. Future systems should prioritize actionable intelligence delivery, ensuring that technical advances translate into improved disaster response capabilities. This includes developing standardized evaluation protocols and establishing partnerships with operational agencies to validate real-world utility.
These research directions collectively aim to transform typological damage assessment from a promising research concept into a robust operational tool for disaster response and recovery planning.
@@ 6 Conclusion
This study introduced DamageCAT, a framework for typology-based building damage assessment that advances beyond traditional severity-focused methods. We contribute the BD-TypoSAT dataset, containing human-annotated satellite imagery with four distinct damage categories (partial roof damage, total roof damage, partial structural collapse, and total structural collapse), addressing the need for descriptive training data that captures the multifaceted nature of structural damage. Our hierarchical U-Net-based transformer architecture analyzes pre- and post-disaster satellite imagery pairs to produce detailed damage classification maps, achieving robust performance with 0.737 IoU and 0.846 F1-score across all damage types.
The framework’s key advancement lies in shifting from broad severity categories to typology-based assessment, providing the detailed information necessary for effective resource allocation. By distinguishing roof damage from structural collapse, emergency responders can prioritize personnel, materials, and financial resources more strategically. The pre-post disaster change detection approach also reduces false positives by filtering out pre-existing building conditions, enhancing reliability for time-critical response situations.
Several noteworthy limitations should be acknowledged. Dataset imbalance, particularly for total structural collapse, reflects real-world damage distributions but indicates the need for expanded, balanced datasets. The reliance on nadir satellite imagery may obscure facade-level or internal damage features that oblique perspectives could reveal. Additionally, training on a single event (Hurricane Ida) in one region (New Orleans, Louisiana) limits generalizability, while the transformer architecture’s computational demands may challenge real-time deployment in resource-constrained environments.
Future research should prioritize balanced data collection across multiple hazards and geographic regions, integration of multi-modal sources (oblique aerial imagery, street-view data, LiDAR), and model optimization for real-time inference. Developing interpretable AI methods could enhance user trust by revealing visual features driving classifications. Cross-regional validation and field testing will be essential for refining the typology-based paradigm and ensuring widespread adoption in operational disaster management contexts.
@@
Data and Code availability The data that support the findings of this study are publicly available at Zenodo, and the code is available at GitHub.
Appendix A Supplementary Figures
(a) Per-Class IoU (Validation)
(b) Per-Class F1-Score (Validation)
Figure 8: Bar charts of (a) per-class IoU and (b) per-class F1-score, derived from validation metrics. For each class, bars represent the highest score achieved during training for each of the four runs, along with the average across runs. Error bars on the average indicate standard deviation.
(a) Per-Class Precision (Validation)
(b) Per-Class Recall (Validation)
Figure 9: Bar charts of (a) per-class Precision and (b) per-class Recall, derived from validation metrics. For each class, bars represent the highest score achieved during training for each of the four runs, along with the average across runs. Error bars on the average indicate standard deviation.
Figure 10: Progression of validation metrics throughout training epochs for Run 1. The plot displays: overall performance (IoU, F1, Precision, Recall), training vs. validation loss, and per-class IoU, F1-score, Precision, and Recall.
Figure 11: Progression of validation metrics throughout training epochs for Run 2. The plot displays: overall performance (IoU, F1, Precision, Recall), training vs. validation loss, and per-class IoU, F1-score, Precision, and Recall.
Figure 12: Progression of validation metrics throughout training epochs for Run 3. The plot displays: overall performance (IoU, F1, Precision, Recall), training vs. validation loss, and per-class IoU, F1-score, Precision, and Recall.
Figure 13: Progression of validation metrics throughout training epochs for Run 4. The plot displays: overall performance (IoU, F1, Precision, Recall), training vs. validation loss, and per-class IoU, F1-score, Precision, and Recall.